A Pillar is dug into the ground 1 third of its length.
half of the length is in water.
2 meters peak out of the water.
How long is the pillar?

Answer by Theo(13342)

A Pillar is dug into the ground 1 third of its length.

this means 1/3 of the length of the pillar is in the ground while 2/3 is above the ground.

half of the length is in water.

you've got 1/3 of the length in the ground and 1/2 of the length in the water.

the total length of the pillar that is either in the ground or in the water is 1/3 + 1/2 = 2/6 + 3/6 = 5/6.

that means that 1/6 of the pillar is neither in the ground or in the water.

let x equal the length of the pillar.

since 2 meters is neither in the ground nor in the water, then 2 meters represents 1/6 of the length of the pillar.

1/6 * x = 2
multiply both sides of this equation by 6 to get x = 12 meters in length total.

1/3 of 12, or 4 meters, is in the ground.
1/2 of 12, or 6 meters, is in the water.
1/6 of 12, or 2 meters, is not in the ground or in the water.